OpenStudy (anonymous):

Solve for y. y-2/y-6 = y+1/y-5+1

OpenStudy (anonymous):

\[\frac{ y-2 }{ y-6 } =\frac{ y+1 }{ y-5+1 }\] ??

OpenStudy (nincompoop):

coocookakabird, how about you try to solve it how do we know what you know if you're not even going to try and just wait for someone to feed you solutions

OpenStudy (anonymous):

I'm sorry but i didn't study math from 2 years i need to review it but today i have to finish something before

OpenStudy (anonymous):

I will try to solve it

OpenStudy (anonymous):

ok i 'll help u

OpenStudy (anonymous):

Thanks my friend ^_^

OpenStudy (anonymous):

\[y-2divy-6=y +1divy -5 +1\]

OpenStudy (anonymous):

1 is not with y-5

OpenStudy (anonymous):

\[\frac{ y-2 }{ y-6 } =\frac{ y+1 }{ y-5 } +1 \] ??

OpenStudy (anonymous):

@kaka2266 is this right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok first u have to add 1 to \[\frac{ y+1 }{ y-5 }\]

OpenStudy (anonymous):

it already have 1 right

OpenStudy (anonymous):

it's

OpenStudy (anonymous):

i think y=7.5

OpenStudy (anonymous):

did any one solve it?

OpenStudy (anonymous):

i think 7.5 is not correct

OpenStudy (anonymous):

Ammm

OpenStudy (anonymous):

i will try it again...did u find it?

OpenStudy (anonymous):

i know how to find it

OpenStudy (anonymous):

Hhhhh how

OpenStudy (anonymous):

do u find \[\frac{ y+1 }{ y-5 } +1\] ?

OpenStudy (anonymous):

\[y + 1 \div y-5 + 1\div1\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

\[\frac{ a }{ b } +\frac{ c }{ d }=\frac{ ad+bc }{ bd }\]

OpenStudy (anonymous):

so it will be just y+1/y-5

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

y=1=y-5/y-5

OpenStudy (anonymous):

y+1+y-5/y-5

OpenStudy (anonymous):

\[\frac{ y+1 }{ y-5} +\frac{ 1 }{ 1 } =\frac{ (y+1) +(y-5)}{y-5 }\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

okay so y-2/y-6=(y=1)=(y-5)/y-5

OpenStudy (anonymous):

y-2/y-6+(y+1)+(y-5)/y-5

OpenStudy (anonymous):

how do u find that term?

OpenStudy (anonymous):

y-2/y-6=(y+1)+(y-5)/y-5 hate that

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think by taking term 2 into the other side?

OpenStudy (anonymous):

but u can simplify the right tem

OpenStudy (anonymous):

i know it will be 2y-4/y-5

OpenStudy (anonymous):

yes u r right

OpenStudy (anonymous):

now i will take it to other side

OpenStudy (anonymous):

now u can use this method \[\frac{ a }{ b }=\frac{ c }{ d }------>ad=bc\]

OpenStudy (anonymous):

ur method is right too

OpenStudy (anonymous):

1 min

OpenStudy (anonymous):

yaah i used urs

OpenStudy (anonymous):

but that is complex

OpenStudy (anonymous):

good

OpenStudy (anonymous):

-y^2+9y-14=0 is that right

OpenStudy (anonymous):

yeah it is right

OpenStudy (anonymous):

(y+7)(y+2)?

OpenStudy (anonymous):

no plz be careful,do it again

OpenStudy (anonymous):

this is ur equation....y^2-9y+14=0

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so (y-7)(y-2)

OpenStudy (anonymous):

y=7,2 right?

OpenStudy (anonymous):

RIGHT

OpenStudy (anonymous):

^_^ thanks a lot for ur time man.

OpenStudy (anonymous):

u can examine ur answers

OpenStudy (anonymous):

NP

OpenStudy (anonymous):

Yup i will